Grade 11Mechanics

PLEASE ANSWER ASAP
A particle is moving with a constant acceleration of a=3i+4j.At time t=0 it starts from rest what is the speed of the particle at time t=1 sec.

MANAS GOEL
8 Years agoGrade 11

1 Answer

hp4evr_sd
ApprovedApproved Tutor Answer8 Years ago
R= root ( 9+16+24cos90°) =root 25 =5
So, a=5m/s^2
Now applying eqns of motion,
a=5m/s^2 , u=0, t = 1 sec , v=?
 
v=u+at
v=0+5×1
v=5m/s
 
So,its speed at t=1 is 5m/s.
Hope it helps.